### v1

# A Path Integral Treatment of Time-dependent Dunkl Quantum Mechanics

Abstract
This paper presents an analytical treatment of the path integral formalism for time-dependent quantum systems within the framework of Wigner-Dunkl mechanics, emphasizing systems with varying masses and time-dependent potentials. By employing generalized canonical transformations, we reformulated the path integral to develop an explicit expression for the propagator. This formalism is applied to specific cases, including a Dunkl-harmonic oscillator with time-dependent mass and frequency. Solutions for the Dunkl-Caldirola-Kanai oscillator and a model with a strongly pulsating mass are derived, providing exact propagator expressions and corresponding wave functions. These findings extend the utility of Dunkl operators in quantum mechanics, offering new insights into the dynamics of time-dependent quantum systems.
1 Introduction
Recently, there has been a significant interest in the exploration of time-dependent quantum systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. This is because these systems have important applications in various fields of physics, including quantum optics [11], cosmology [12], nanotechnologies [13], and plasma physics [14]. The diverse applications of these systems have prompted the development of various methods and techniques to study such systems. Notable among these are the Lewis-Riesenfeld invariant operator method [15, 16], path integral formalism [17], and coherent states [18]. The Lewis-Riesenfeld invariant operator method is an approach that enables the determination of a complete set of solutions to the Schrödinger equation for time-dependent potentials by utilizing the eigenstates of an invariant operator. This method consists of two main steps: first, constructing the invariant operator, and second, expressing Schrödinger’s wave function using the eigenstates of this invariant operator, combined with a time-dependent phase factor. On the other hand, the Feynman path integral formalism provides an approach to solving quantum mechanical problems [19] alternative to the well-known methods of Heisenberg and Schrödinger. However, this formulation’s applicability has been limited because explicit expressions for propagators are available only for a few cases. In the path integral approach, various time-dependent quantum systems have been studied, such as the harmonic oscillator with varying frequency or mass [10, 17, 20, 21, 22, 23, 24, 25], the infinite potential well [17], and the time-dependent model of the form [26].
In 1950, Wigner studied the relationship between quantum dynamics and commutation relations. He demonstrated that the idea of quantization, represented by the canonical commutation relation is a solution to the Heisenberg equation of motion, but it is not the only possible solution [27]. Just a year later, Yang considered the problem that had been addressed by Wigner [28]. By using a more precise definition of Hilbert space and introducing another condition involving a rigorous theorem about expanding mathematical objects, Yang showed that commutation relations could be deduced uniquely if a reflection operator, , is taken into consideration:
where denotes the Wigner parameter. In 1989, Dunkl introduced a new type of differential operator , defined as a linear combination of partial derivatives weighted by the reflection operator, allowing the study of special functions and their properties in a more general setting [29, 30]. The one-dimensional version of the Dunkl operator is given by
Quantum mechanical systems involving the reflection operator have recently found applications in physics related, in particular, to anyons [31, 32], supersymmetry [33, 34, 35, 36, 37, 38, 39], integrable systems [33, 40], non-commutative geometry [32], and PT-symmetry [41, 42]. Substituting the ordinary partial derivative with the Dunkl derivative has found applications in a variety of physics fields, enabling the exact study of different quantum mechanics problems using various methods [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Recently, four papers have explored the development of path integral formalism within the context of Wigner-Dunkl quantum mechanics [64, 65, 66, 67]. One of these papers focused on constructing the propagator for one-dimensional quantum systems incorporating the Dunkl derivative [64]. The authors employed the configuration space representation and Feynman’s path integral approach in their analysis. Their findings revealed that the Dunkl path integral kernel can be expressed as a combination of even and odd kernel propagators.
Our aim in this paper is to explore the Dunkl path integral formalism for the propagator in the presence of a time-dependent potential. The treatment is mainly based on suitably chosen generalized canonical transformations and auxiliary equations. This paper is organized as follows: In Sec. 2, we reformulate the path integral formalism for the propagator of the time-dependent system. In Sec. 3, we consider the case of the Dunkl-harmonic oscillator with a time-dependent mass and frequency, and we calculate the relative propagator and the wave functions. In Sec. 4, we illustrate the general procedure by considering two models of varying mass. Finally, Sec. 5 contains the conclusion.
2 Time-dependent propagator
We start by outlining a method to compute the path integral for non-relativistic Wigner-Dunkl quantum systems with a time-dependent mass in the presence of general time-dependent potentials. This approach employs explicitly time-dependent spacetime transformations to facilitate the calculations, with the systeme’s dynamics governed by a Wigner-Dunkl Hamiltonian:
where
In the functional integral formalism, the Wigner-Dunkl-propagator corresponding to the system described by Eq. (3) can be expressed as [64, 65],
where denotes the sign function, represents the even kernel
and corresponds to the odd kernel
propagators. To keep the mass term in the kinetic energy constant, we can convert the propagator into an appropriate form by continuing the treatment in phase space as described in [17]. In this context, the kernel propagator in phase space takes the form:
where
and
Here, takes the value for the even case and for the odd case.
To enhance the manageability of the path integral and facilitate its calculations, it is useful to utilize the method of canonical transformations [17], specifically by performing a time-dependent canonical transformation defined as,
with the generating function
which can be derived from the following equations of classical mechanics
As a consequence of these transformations, we reformulate the action defined in Eq. (8) in terms of the new conjugate variables as follows:
where
is the new Hamiltonian which governs the movement in the system , and
Here, the dot notation over the variable refers to the time derivative. The transformation subsequently alters the integral measure; following the procedures described in [17], we find that the path integral measure transforms as follows:
From Eqs. (15) and (18), it follows that the evolution of the physical system in the new coordinate system is characterized by
We now utilize a transformation of the time variable, changing to :
so that becomes
where , . At this stage, we notice the presence of an additional term in the propagator expression given in Eq. (21). To eliminate this term, we introduce the following canonical transformation :
with the generating function given by;
In this case, the motion of the particle within the system is governed by the new Hamiltonian:
where
Thus, the propagator corresponding to the system in Eq. (24) can be expressed as:
We now decompose the kernel propagator into two parity components, even and odd, distinguished by their respective eigenvalues, . The propagator for the even parity case is given by:
and the odd parity case propagator is:
3 Harmonic oscillator
The result presented in Eq. (26) is significant and notably interesting due to its relative simplicity. In this section, we demonstrate how this result can be applied, highlighting a noteworthy application of substantial epistemological relevance: the harmonic oscillator with time-dependent mass and frequency,
Employing the potential in Eq. (26), the propagator takes the form:
We now impose a constraint on by setting the global time-dependent frequency in Eq. (30) to a constant,
Consequently, Eq.(24) becomes
which is referred to as the auxiliary equation. By substituting Eq. (31) into Eq. (30), we obtain the final expression for the propagator of the time-dependent Dunkl-harmonic oscillator as:
where is the modified Bessel function. Returning to the original variable and applying the decomposition outlined in Eq. (5), we can express the even and odd kernel propagators as follows:
where
Setting
in the expression for the kernel propagators given in Eq. (33), and employing the Hille-Hardy formula [68],
we derive the even and odd spectral decompositions of kernel propagators:
where
is the even wave functions and
the odd wave functions, are the associated Legendre or Laguerre polynomial. It is noteworthy that the eigenfunctions we obtained coincide with those derived in [64] through the Lewis-Riesenfeld method. Then, with the help of the deformed exponential function [68], defined as:
we express the propagator of the time-dependent Dunkl-harmonic oscillator as follows:
where It is important to note that for the case where the mass and frequency are constants, i.e. and , we recover the propagator of the one-dimensional harmonic oscillator within the framework of Wigner-Dunkl quantum mechanics [65],
4 Applications
At this stage, it is necessary to examine specific models of time-varying harmonic oscillators to validate the analytical result presented in Eq. (33). These models should provide well-defined expressions for and that yield exact solutions. In the following subsection, we explore two examples: the Caldirola-Kanai oscillator and a harmonic oscillator with a strongly pulsating mass.
4.1 Caldirola-Kanai Oscillator
Let us begin by examining a harmonic oscillator with a constant frequency and a mass that varies with time in the form of
where is a real number, and is the constant damping coefficient. In this case, the Dunkl-Caldirola-Kanai Hamiltonian for this system can be expressed as:
and the solution of the auxiliary equation, given in Eq. (32), takes the form
where is the reduce frequency
By employing a time transformation defined in Eq. (20), we obtain an exact expression for the propagator as follows:
By inserting this propagator in Eq. (38), we readily derive the wave functions. The even and odd wave functions are expressed, respectively, in terms of Legendre polynomials as follows:
We note that this result is consistent with that obtained in [64].
4.2 The harmonic oscillator with strongly pulsating mass
The harmonic oscillator with a strongly pulsating mass can be used to model the interaction between the electromagnetic field in a Fabry-Perot cavity and a reservoir of resonant two-level atoms. The periodic emission and reabsorption of photons can be described by an oscillator with energy that fluctuates periodically. This system can be represented by an oscillator with a periodically fluctuating energy, which corresponds to a mass that varies periodically over time as,
where is the constant, represents the frequency of mass. In this case, the Hamiltonian transforms into
By inserting the time-dependent mass into the auxiliary Eq.(32), we obtain
where , is the augmented frequency. By inserting Eqs. (51) and (52) into the general propagator, we can derive the time-dependent propagator for a Dunkl-harmonic oscillator with a strongly pulsating mass as follows,
This propagator can be simplified by setting , reducing the result to the propagator of the stationary Dunkl harmonic oscillator [64].
5 Conclusion
Time-dependent quantum mechanics has become a pivotal area of research due to its applications across diverse fields, including plasma physics, quantum optics, and quantum chemistry. While the path integral method is widely used in time-dependent quantum systems, its application within the Dunkl formalism remains largely unexplored. In this study, we extend the path integral approach to time-dependent Dunkl quantum mechanics by reducing it to a stationary form through explicit generalized canonical and time transformations.
More precisely, we derived a general expression for the propagator and developed explicit formulations for the Dunkl harmonic oscillator, with a particular focus on two analytically solvable cases involving harmonic oscillators with time-dependent mass. This framework was applied to the Dunkl-harmonic oscillator with time-dependent mass and frequency, enabling the calculation of exact propagators and corresponding wave functions. These results underscore the applicability of Dunkl operators within time-dependent quantum systems, highlighting their versatility in addressing non-stationary dynamics. Subsequently, the effectiveness of this approach was further validated by its application to two illustrative models: the Dunkl-Caldirola-Kanai oscillator, characterized by an exponentially varying mass, and a harmonic oscillator with a strongly pulsating mass. These examples provided concrete instances of how Dunkl mechanics can effectively accommodate complex dependencies of mass and frequency, thereby expanding the toolkit for analyzing time-dependent quantum systems.
Overall, this study demonstrates the capacity of the Dunkl formalism to address complex time-dependent quantum problems and derive exact solutions in cases that pose significant analytical challenges. This approach broadens the scope of Dunkl mechanics for applications in both theoretical and applied physics, offering novel insights into systems where time-dependent characteristics and reflection symmetries play a critical role.
Acknowledgments
B. C. L. is grateful to Excellence Project PřF UHK 2211/2023-2024 for the financial support.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Conflict of Interests
The authors have no conflicts to disclose.